# Learning holographic QCD with unflavoured meson spectra

### v1

@@ Appendix A Numerical Routine
where $B_{1}^{\prime}(z)=\partial_{z}B_{1}(z)$, $\pi_{i}=\pi(z_{i})$ and $\xi_{i}=\xi(z_{i})$. For the boundary condition, $\pi_{0}=\pi_{N}=0$ and $\xi_{0}=\xi_{N}=0$, we can write this equation in terms of a matrix as
@@
$M_{1}$ $=\begin{pmatrix}\frac{-2}{h^{2}}&\left(\frac{1}{h^{2}}+\frac{\omega(z_{1})}{2h}\right)&0&\cdots&0\\ \left(\frac{1}{h^{2}}-\frac{\omega(z_{2})}{2h}\right)&\frac{-2}{h^{2}}&\left(\frac{1}{h^{2}}+\frac{\omega(z_{2})}{2h}\right)&\cdots&0\\ 0&\left(\frac{1}{h^{2}}-\frac{\omega(z_{3})}{2h}\right)&\frac{-2}{h^{2}}&\cdots&0\\ \vdots&\vdots&\vdots&\ddots&\vdots\\ 0&0&0&\cdots&\frac{-2}{h^{2}}\end{pmatrix}$, (45)
@@
$M_{3}$ $=\begin{pmatrix}\left(\frac{-2}{h^{2}}-C(z_{1})\right)&\left(\frac{1}{h^{2}}-\frac{B_{1}^{\prime}(z_{1})}{2h}\right)&0&\cdots&0\\ \left(\frac{1}{h^{2}}+\frac{B_{1}^{\prime}(z_{2})}{2h}\right)&\left(\frac{-2}{h^{2}}-C(z_{2})\right)&\left(\frac{1}{h^{2}}-\frac{B_{1}^{\prime}(z_{2})}{2h}\right)&\cdots&0\\ 0&\left(\frac{1}{h^{2}}+\frac{B_{1}^{\prime}(z_{3})}{2h}\right)&\left(\frac{-2}{h^{2}}-C(z_{3})\right)&\cdots&0\\ \vdots&\vdots&\vdots&\ddots&\vdots\\ 0&0&0&\cdots&\left(\frac{-2}{h^{2}}-C(z_{N-1})\right)\end{pmatrix}$ (47)
We solve the generalized eigenvalue equation given in Eq. 43 using scipy.linalg.eigvals giving us the non-dimensional mass eigenvalues $m_{n}^{2}$. The eigenvalues obtained are sorted in ascending order, and the first few eigenvalues are taken as the mass spectrum of the pseudo-scalar mesons.
@@ Appendix B Training details
In this section, we provide the details of the training procedure and various hyperparameters used. We have taken $z_{max}=10$ and $h=0.1$, which gives $N=100$ discretization points. The $a$ and $b$ values in Eq. 24 are taken as 0.01 GeV3 and 0.03 GeV3 respectively.
@@
where, $E_{\mu,n}$ is the error for the $n^{th}$ mass of the meson as given in Eq. 30. Note that $\Lambda_{\mu,n}$ needs to be detached from the computational graph while calculating the loss to avoid any unwanted gradients. Additionally we have chosen $\Lambda_{v^{\prime}}=10$, $\Lambda_{A}=1$ and $\Lambda_{\text{pot}}=1$ for the other loss components given in Eq. 36.
@@
We define a search space for the initial value of $k_{1}$ and $k_{2}$ as [15,15][-15,15] and run the optimization for 100 trials. In each trial, we train the model for 15,000 iterations. The best values of $k_{1}$ and $k_{2}$ obtained from the optimization are $k_{1}=-7.4$ and $k_{2}=12.2$. These values are used as the initial guess for $k_{1}$ and $k_{2}$ in the subsequent training runs.
@@
For the training, we have used two Adam optimizers [52] with different learning rates. The first Adam optimizer is used to update the weights and biases of the neural network, with an initial learning rate of $5\times 10^{-4}$, which is reduced in further iterations using a learning rate scheduler. The learning rate scheduler reduces the learning rate by a factor of 0.95 if the loss does not decrease for 100 iterations. The minimum learning rate is set to $10^{-8}$ for this optimizer. The second Adam optimizer is used to train the model parameters $k_{1}$, $k_{2}$, $L$, and $\theta$, with an initial learning rate of $5\times 10^{-2}$, along with a learning rate scheduler that reduces the learning rate by a factor of 0.6 if the loss does not decrease for 200 iterations. The minimum learning rate is set to $10^{-5}$ for this optimizer.
We have also implemented gradient clipping for all the parameters with a maximum norm of 1.0. Additionally, we have used double precision (float64) for better numerical stability. We train the model for 15,000 iterations.
@@
1 $8.59\times 10^{-3}$ 0.274 7.39-7.39 $16.26$ $1.49$ $2.97$
2 $9.56\times 10^{-3}$ 0.567 7.19-7.19 $15.44$ $1.83$ $1.40$
3 $8.24\times 10^{-3}$ 1.323 7.34-7.34 $16.10$ $1.70$ 0.20-0.20
4 $7.90\times 10^{-3}$ 0.374 7.00-7.00 $14.58$ $1.54$ $2.67$
5 $6.56\times 10^{-3}$ 0.431 7.50-7.50 $15.47$ $1.57$ $2.55$
6 $8.72\times 10^{-3}$ 0.268 6.30-6.30 $13.55$ $1.61$ $4.02$
7 $7.61\times 10^{-3}$ 0.610 7.55-7.55 $16.82$ $1.59$ $0.91$
8 $8.82\times 10^{-3}$ 0.359 8.11-8.11 $18.72$ $1.53$ $2.49$
9 $\mathbf{5.00\times 10^{-3}}$ 0.455 9.85-9.85 $21.28$ $1.75$ $4.42$
10 $9.53\times 10^{-3}$ 0.833 9.51-9.51 $22.52$ $2.44$ $2.85$
Table 5: Comparison of the minimum training loss for $\mathcal{L}_{\text{mass}}^{(\rho,a_{1},a_{2},f_{0})}$ and the $\pi$ mass loss $\mathcal{L}_{\text{mass}}^{(\pi)}$ across all runs, along with the corresponding model parameters $(k_{1},k_{2},L,\theta)$. Run 9 achieved the lowest training loss $\mathcal{L}_{\text{mass}}^{(\rho,a_{1},a_{2},f_{0})}$ (This loss is weighted according to their appropriate $\Lambda$ values given by the Eq. 49), while Run 6 yielded the lowest $\pi$ mass loss $\mathcal{L}_{\text{mass}}^{(\pi)}$ (This loss is unweighted i.e. $\Lambda_{\pi,n}=1$), despite not being trained for it.
We run the training multiple times and choose only those runs where the minimum total loss converges to a value below $10^{-2}$. The loss curve for a representative run is shown in Fig. 6. The training loss for each run is given in Table 5. We trained the model on an Intel Xeon W-2245 CPU (3.90 GHz, 16 GB RAM), which took less than 30 minutes per run.

### v2

@@ Appendix A Numerical Routine
where $B_{1}^{\prime}(z)=\partial_{z}B_{1}(z)$, $\pi_{i}=\pi(z_{i})$ and $\xi_{i}=\xi(z_{i})$. From the Dirichlet boundary condition at $z=0$, we have $\pi_{0}=0$ and $\xi_{0}=0$. From the Neumann boundary condition $\partial_{z}\pi=\partial_{z}\xi=0$ at $z=z_{max}$, we have $\pi_{N}=\pi_{N-1}$ and $\xi_{N}=\xi_{N-1}$. Substituting these boundary conditions in the above equation, we get
@@
$M_{1}$ $=\begin{pmatrix}\frac{-2}{h^{2}}&\left(\frac{1}{h^{2}}+\frac{\omega(z_{1})}{2h}\right)&0&\cdots&0\\ \left(\frac{1}{h^{2}}-\frac{\omega(z_{2})}{2h}\right)&\frac{-2}{h^{2}}&\left(\frac{1}{h^{2}}+\frac{\omega(z_{2})}{2h}\right)&\cdots&0\\ 0&\left(\frac{1}{h^{2}}-\frac{\omega(z_{3})}{2h}\right)&\frac{-2}{h^{2}}&\cdots&0\\ \vdots&\vdots&\vdots&\ddots&\vdots\\ 0&0&0&\cdots&\frac{-2}{h^{2}}+\frac{1}{h^{2}}+\frac{\omega(z_{N-1})}{2h}\end{pmatrix}$, (45)
@@
$M_{3}$ $=\begin{pmatrix}\frac{-2}{h^{2}}-C(z_{1})&\frac{1}{h^{2}}-\frac{B_{1}^{\prime}(z_{1})}{2h}&0&\cdots&0\\ \frac{1}{h^{2}}+\frac{B_{1}^{\prime}(z_{2})}{2h}&\frac{-2}{h^{2}}-C(z_{2})&\frac{1}{h^{2}}-\frac{B_{1}^{\prime}(z_{2})}{2h}&\cdots&0\\ 0&\frac{1}{h^{2}}+\frac{B_{1}^{\prime}(z_{3})}{2h}&\frac{-2}{h^{2}}-C(z_{3})&\cdots&0\\ \vdots&\vdots&\vdots&\ddots&\vdots\\ 0&0&0&\cdots&\frac{-2}{h^{2}}-C(z_{N-1})+\left(\frac{1}{h^{2}}+\frac{-B_{1}^{\prime}(z_{N-1})}{2h}\right)\end{pmatrix}$ (47)
We solve the generalized eigenvalue equation given in Eq. 43 using scipy.linalg.eigvals giving us the non-dimensional mass eigenvalues $m_{n}^{2}$. The eigenvalues obtained are sorted in ascending order, and the first few eigenvalues are taken as the mass spectrum of the pseudo-scalar mesons. We have observed that unlike the other mesons, the ground state of the $\pi$ meson needs a finer discretization to be accurately captured, which is why, for this case we have chosen $h=0.02$.
@@ Appendix B Training details
In this section, we provide the details of the training procedure and various hyperparameters used. We have taken $z_{max}=15$ and $h=0.15$, which gives $N=100$ discretization points. The $a$ and $b$ values in Eq. 24 are taken as 0.01 GeV3 and 0.03 GeV3 respectively.
@@
where, $E_{\mu,n}$ is the error for the $n^{th}$ mass of the meson as given in Eq. 30. While calculating $\Lambda_{f_{0},1}$, we scale the error $E_{f_{0},1}$ by a factor of 0.1 to give it less weightage in the loss function, since the $f_{0}(500)$ meson is a broad resonance and its mass is not well defined. Note that $\Lambda_{\mu,n}$ needs to be detached from the computational graph while calculating the loss to avoid any unwanted gradients. Additionally we have chosen $\Lambda_{v^{\prime}}=10$, $\Lambda_{A}=1$ and $\Lambda_{\text{pot}}=1$ for the other loss components given in Eq. 36.
@@
We define a search space for the initial value of $k_{1}$ and $k_{2}$ as [20,20][-20,20] and run the optimization for 100 trials. In each trial, we train the model for 10,000 iterations. The best initial guess values of $k_{1}$ and $k_{2}$ obtained from the optimization are $k_{1}=-2.2$ and $k_{2}=12.0$. These values are used as the initial guess for $k_{1}$ and $k_{2}$ in the subsequent training runs.
@@
For the training, we have used two Adam optimizers [46] with different learning rates. The first Adam optimizer is used to update the weights and biases of the neural network, with an initial learning rate of $1\times 10^{-3}$, which is reduced in further iterations using a learning rate scheduler. The learning rate scheduler reduces the learning rate by a factor of 0.95 if the loss does not decrease for 100 iterations. The minimum learning rate is set to $10^{-8}$ for this optimizer. The second Adam optimizer is used to train the model parameters $k_{1}$, $k_{2}$, $L$, and $\theta$, with an initial learning rate of $5\times 10^{-2}$, along with a learning rate scheduler that reduces the learning rate by a factor of 0.6 if the loss does not decrease for 200 iterations. The minimum learning rate is set to $10^{-5}$ for this optimizer.
We have also implemented gradient clipping for all the parameters with a maximum norm of 1.0. Additionally, we have used double precision (float64) for better numerical stability. We train the model for 10,000 iterations.
@@
1 $5.35\times 10^{-3}$ 0.378 4.01-4.01 $7.45$ $1.23$ $4.53$
2 $6.42\times 10^{-3}$ 0.359 4.84-4.84 $9.59$ $1.20$ $5.14$
3 $7.82\times 10^{-3}$ 0.586 3.51-3.51 $6.62$ $1.60$ $5.11$
4 $7.38\times 10^{-3}$ 0.400 2.93-2.93 $6.04$ $1.00$ $1.57$
5 $6.76\times 10^{-3}$ 0.307 2.54-2.54 $4.40$ $1.00$ $4.10$
6 $8.16\times 10^{-3}$ 0.575 4.15-4.15 $7.98$ $1.61$ $4.63$
7 $\mathbf{4.44\times 10^{-3}}$ 0.510 5.13-5.13 $10.16$ $1.39$ $4.56$
8 $4.94\times 10^{-3}$ 0.305 5.24-5.24 $10.75$ $1.00$ $3.33$
9 $6.89\times 10^{-3}$ 0.474 5.29-5.29 $10.87$ $1.31$ $2.12$
10 $4.72\times 10^{-3}$ 0.353 5.68-5.68 $12.13$ $1.00$ $1.88$
Table 5: Comparison of the minimum training loss for $\mathcal{L}_{\text{mass}}^{(\rho,a_{1},a_{2},f_{0})}$ and the $\pi$ mass loss $\mathcal{L}_{\text{mass}}^{(\pi)}$ across all runs, along with the corresponding model parameters $(k_{1},k_{2},L,\theta)$. Run 7 achieved the lowest training loss $\mathcal{L}_{\text{mass}}^{(\rho,a_{1},a_{2},f_{0})}$, while Run 8 yielded the lowest $\pi$ mass loss $\mathcal{L}_{\text{mass}}^{(\pi)}$, despite not being trained for it.
We run the training multiple times and choose only those runs where the minimum total loss converges to a value below $10^{-2}$. The loss curve for a representative run is shown in Fig. 7. The training loss for each run is given in Table 5. We trained the model on an Intel Xeon W-2245 CPU (3.90 GHz, 16 GB RAM), which around 20 minutes per run.